# Exploring the Nostr Ecosystem: A Study of Decentralization and Resilience

### v1

@@ 3.2. Event Dataset
Event collection.  We connect to each relay in our dataset to gather all posts, users and zap related events. We divide the collection process into several parallelized jobs and distributes the workload evenly across the relays, preventing any single relay from being overwhelmed.
Text Notes (Posts).  A text note in Nostr is a kind-1 event, which is the fundamental kind of event, used for user posts. This is roughly equivalent to a tweet on Twitter. However, in contrast to Twitter, we emphasize that an event can be replicated across multiple relays (but is identified by a unique ID). Thus, we regard each unique ID as an individual event and record the relays in which it is replicated. We crawl 712 relays for the time period from 2023-07-01 to 2023-12-31, and get a total of 17,851,871 text notes.
Users.  Each unique pubkey represents an individual user. We follow this convention throughout the remainder of the paper. We identify active users who have posted at least once between 2023-07-01 and 2023-12-31 from our text note dataset. We obtain 1,558,891 unique pubkeys from the text note dataset. Subsequently, we construct the user dataset by querying the relays for user profiles (kind-0 event) and the user’s following list (kind-3 event) using the obtained public keys.
Zaps.  To investigate the financial incentives of operating a relay, we collect kind-9735 events associated with the account of relay operators, which is used for the zap transactions as described in §2. We crawl the relays for the time period from 2023-07-01 to 2023-12-31, and get a total of 252,905 zap transactions, with the amount summed up to be 305,427,302 Satoshi (1 Bitcoin = 100 million Satoshi). At the time of writing, it is roughly 137,000 USD.
@@ 4. Characterizing Nostr Relays
Nostr is designed to offer high-availability decentralized support for content sharing. Thus, a key design goal is to decentralize operation away from single points of failure, and maintain high availability. The success of this can be measured through several metrics, including the distribution of posts and users across relays (to identify any points of failure), alongside the hosting patterns of the relays themselves. In this section, we focus on studying the properties of the Nostr relays and explore to what extent this relay paradigm achieves high level of decentralization while maintaining high availability.
@@ 4.1. Distribution of Posts & Users
Figure 2. (a) The distribution and CDF of the number of posts in a relay; (b) The distribution and CDF of the number of users in a relay.
We first evaluate the degree of decentralization within Nostr at the relay level. To do this, we inspect the distribution of posts and users across the relays, in terms of the number of posts in the relay and the number of users using the relay. We define that a user is using a relay if they have at least one post in the relay. Previous research has highlighted user-centralization and content-centralization problems in the Fediverse (Raman et al., 2019). That is, a small number of instances host a vast majorly of posts and users. While we aim to determine if these findings also apply to Nostr, we also note a vital difference: in Nostr, a user and a post can exist in multiple relays. We posit that this uniqueness may result in distinctive characteristics.
Distribution of Posts & Users.  Figure 2a illustrates the distribution and cumulative distribution function (CDF) of the number of posts in a relay. We observe that the top relay hosts 73% of the posts. At first glance, this may suggest a highly centralized distribution. However, it is important to consider that a post can exist in multiple relays. Taking this into account, the result reveals that although 178 relays (25%) host a significant proportion (more than 5.6%) of the posts, the sum of the proportions is far larger than 100%, reaching 3200%. This indicates that posts are extensively replicated and distributed across multiple relays. This could be viewed as a decentralized backup of posts, serving to mitigate the risk of failure or censorship in a single relay. Therefore, the result stands as a favorable signal of decentralization.
Figure 2b illustrates the distribution and CDF of the number of users in a relay. We observe a similar distribution to the distribution of posts. We find that the proportions across all relays summed up to be 1310%. This suggests that users are actively utilizing various relays for publishing (and likely retrieving) posts. Once again, this highlights Nostr’s distinctive feature of detaching users from specific servers, potentially mitigating the risk of user-centralization.
Overall, the distribution patterns of posts and users demonstrates the effective decentralization of Nostr at the relay level. And, unlike Fediverse, we do not find evidence for excessive user-centralization or post-centralization at the relay level in Nostr.
@@ 4.2. Relay Hosting
In this section, we evaluate the degree of decentralization of Nostr at country and autonomous systems (AS) level. Like instances in the Fediverse, relays in Nostr follow a bottom-up approach, where administrators have the autonomy to choose where to host their relays. Prior work revealed that there is infrastructure-centralization in Fediverse (Raman et al., 2019). That is, instances and users and posts in the instances are centralized in a few countries and Autonomous Systems (ASes). Thus, we hypothesize that Nostr may encounter similar challenges whereby the majority of relays fall under the remit of a small number of countries or ASes. To test this, we investigate the hosting of relays in Nostr to see if relays, posts or users are overly centralized in certain locations.
Countries.  Figure 3a displays the CDF of the percentage of relays, hosted posts, and users in different countries. Additionally, Figure 3c illustrates the percentage of relays, hosted posts, and users for the top 15 countries ranked by the number of relays.
In stark contrast to the Fediverse, we observe that the distribution of relays is quite decentralized, with around 52% of countries having fewer than two relays. That said, upon closer inspection, we see that users and posts are concentrated in the top countries, such as the US, which hosts 85% of posts and 30% of users. Nonetheless, it is crucial to recall that, unlike Fediverse instances, a post and a user can belong to multiple relays, which creates backups that will persist even if some relays fail. As a result, most of the top 15 countries host at least 20% of the posts, and Canada, France, Germany, and Finland host more than 40%. This distribution is actually a positive indicator of decentralization, as it suggests that a significant portion of Nostr posts and users are spread across different countries. This outcome appears to be more favorable compared to Mastodon, where a study in 2019 revealed that 89.1% of all toots were concentrated on instances in Japan, the US, and France (Raman et al., 2019).
ASes.  Figure 3b displays the CDF of the percentage of relays, hosted posts, and users in different ASes. Additionally, Figure 3d illustrates the percentage of relays, hosted posts, and users for the top 15 ASes ranked by the number of relays.
We observe a similar trend to countries, as the distribution of relays across ASes is also surprisingly decentralized: 64% of ASes host only one relay. As shown in Figure 3d, among the top 15 ASes, all of them host more than 15% of the posts. Once again, this serves as a positive indicator of decentralization. That said, there are certain hot spots. We note that certain relays utilize Content Delivery Networks (CDNs), with Cloudflare and Akamai hosting around 80% of the posts respectively. We note that such CDNs do not host the relay. Instead, they support its object delivery. We are unable to identify the specific servers behind these CDNs, as there is currently no reliable method to do so.
This distribution also contrasts significantly with the findings in the Fediverse. According to the study in 2019 (Raman et al., 2019), toots and users in Mastodon were primarily centralized around three major ASes, as administrators were naturally drawn to well-known and cost-effective cloud providers. One possible explanation for this disparity is that Nostr relays are typically more lightweight than Fediverse instances, requiring fewer computational resources. Consequently, administrators may find better pricing options from smaller providers rather than relying on major providers like AWS.
@@
Figure 3. (a) CDF of the percentage of relays, posts, and users in the countries; (b) CDF of the percentage of relays, posts, and users in the ASes; (c) Percentage of relays, posts, and users in the top 15 regions, ranked by number of relays; (d) Percentage of relays, posts, and users in the top 15 ASes, ranked by number of relays.
@@ 4.3. Relay Availability
A supposed outcome of greater decentralization is superior resilience and availability. Thus, we next explore the availability properties of the relays. In this context, availability refers to the ability for us to establish a connection with the relay. Similar to Fediverse instances, we posit that the voluntary nature of certain relay operations might lead to unique availability characteristics.
Relay Availability.  Based on the relay dataset described in Section 3, we calculate the downtime of the relays from 2023-10-01 to 2023-12-31. The CDF of the downtime percentage is presented in Figure 4a. From the figure, we observe that the availability of relays is relatively good, with 50% of the relays being available for more than 99% of the specified time period. However, there are also some relays with extremely poor availability. Approximately 20% of the relays experience downtime for more than 40% of the time.
However, further investigation reveals that most of these relays are offline for extended periods and do not even reactivate by the end of our measurements. We therefore define relays that have been offline for more than a week and remain so at the conclusion of our measurements as dead. In total, we identify 132 dead relays. When we remove these dead relays from the analysis, as shown by the red line in Figure 4a, the availability of improves significantly, with 90% of the relays being operational for more than 80% of the time.
Overall, this finding confirms that failures are relatively common in Nostr. This broadly aligns with that observed in the Fediverse (Raman et al., 2019) (although the most significant failures can be attributed to dead relays). However, fortunately, in Nostr, users have the flexibility to switch between relays easily. As a result, the impact of relay unavailability in Nostr may not be as severe as in instance unavailability in Fediverse.
Outage Duration.  We next investigate the outage duration. For each outage, we compute its duration and plot the CDF in Figure 4b and 4c . We do not exclude the dead relays in this analysis, as dead relays can be considered as relays with a long outage duration. First, we compare free and paid relays to determine if paid relays also exhibit better performance in terms of shorter outage duration. Interestingly, we observe that the outage durations for paid and free relays are nearly identical, as depicted in Figure 4b. The majority of outages (56%) last for less than 15 minutes. However, there is also a long tail, with 10% of the outages lasting for more than 100 minutes.
Next, we group the relays based on the number of posts to observe if the scale of the relays correlates with the outage duration. As shown in Figure 4c, we notice that larger relays perform better, particularly in the tail end. For the group of relays with over 1 million posts, only 1% of the outages last for more than 1000 minutes, whereas the corresponding figure for the group of relays with less than 100 posts is 88%. This finding suggests that although the downtime may be similar, the outage durations for larger relays are shorter. This implies that while the availability of larger relays may not be consistently stable, it is unlikely for them to experience prolonged outages lasting several hours.
@@
Figure 4. (a) CDF of the percentage of downtime for free, paid, alive, and all relays; (b) CDF of the outage duration for free, paid, and all relays; (c) CDF of the outage duration for relays with different number of posts.
@@ 4.4. Relay Incomes
Figure 5. (a) CDF of the amount of zap and count of zap transactions and zappers for relays; (b) Number of users vs. the estimated 6-month income of free and paid relays. The green line is the estimated 6-month cost of running a relay.
@@
One of the significant challenges within the Fediverse is the cost of running instances, coupled with a lack of incentives for doing so (Allen et al., 2023). This is also one reason why some instances have low availability, with some even being abandoned and subsequently closed. This issue also applies to Nostr, as the lack of financial income may lead to a poor availability of certain relays and the system may become centralized around resourced individuals (who do not require financial support). Therefore, we examine the possible incomes of the relays, and importantly, whether the income generated is sufficient to cover operational costs, and thus to keep the decentralized operation of the Nostr ecosystem. We argue that this is key to relay availability, as a lack of money could result in relays shutting down.
@@
Paid Relays.  One direct source of revenue is the admission fee for paid relays. This fee is a one-time cost for users to publish to the relay. We find that the admission fee is relatively low for some of the relays, with 35% charging less than 1k Satoshi (approximately 0.45 USD). We argue that this minimal fee makes it challenging to sustain the relay’s operation. This implies that, for these relays, the primary purpose of the admission fee is to act as a barrier, allowing only genuine users while keeping spammers out, rather than generating profits. On the other hand, there are also relays that impose a higher admission fee, with over 15% of the relays charging more than 10k Satoshi (approximately \$45 USD). This higher fee may potentially help in covering the expenses associated with operating the relay. We further explore this in the subsequent paragraphs.
@@
Zaps.  Another available revenue stream is through zaps sent by users to relay administrators. To estimate this, we analyze the zap transactions to the public keys specified in relay information. If no public key is provided in the relay information, we consider the zap income as zero. Figure 5a shows the CDF of the total zap amount (We note that 52% of the transactions do not disclose the zap amount publicly), count of zap transactions, and zappers. We find that the majority (64%) of relays do not receive any zaps. Among the remaining relays, some of them do receive a relatively high number of zap transactions, with the 90th percentile at 471. However, the zap amount itself is not substantial, with the 90th percentile receiving a total of 150k Satoshi (approximately 67 USD).
@@
Enough to Cover the Cost?  We finally we assess whether the income generated is sufficient to cover operational costs. For this, we estimate the 6-month income (2023-07-01 to 2023-12-31) of relays. To achieve this, we first calculate the zap amount for the relays. Then, for paid relays, we add the product of the admission fee and the number of active users. For relays operational for less than 6 months, we calculate the daily average income and multiply it by 184 to estimate the income for the 6-month period. Figure 5b illustrates the estimated income vs. the number of users as a representation of the relay’s scale (See Appendix A.2 for figures with different Bitcoin prices). The green line roughly estimates the cost of operating a relay based on the cost of running a Mastodon instance (Mastohost, 2023).
@@
It is evident that, barring a few outliers, most of the free relays with more than 10 users are unable to cover their operating costs. However, most paid relays of medium scale (with 50 - 1k users) are able to cover their operational costs. This constitutes a key threat to long-term viability. It is therefore necessary for relays to obtain additional revenue (often donated by the voluntary administrator). Recall, 52% of the transactions do not disclose the zap amount publicly. However, even if we assume that the actual amount is twice as much, it still remains insufficient.
This result suggests that, akin to Fediverse, most Nostr relays are operated by enthusiasts without clear incentives. The number of dead relays and the lower availability of free relays revealed in §4.3 may be a consequence of this situation. Although a relay’s role in Nostr may not be as critical as an instance in Fediverse, we argue that the lack of clear incentives to operate relays remains a problematic issue.
@@
inline,color=whiteinline,color=whitetodo: inline,color=whiteTakehomes: (i) Nostr achieves greater decentralization than previous Fediverse applications, with posts broadly spread and replicated across multiple relays in 44 countries and 151 autonomous systems. (ii) Unstable availability of relays is still a challenge for Nostr, akin to Fediverse applications. (iii) Financial sustainability may be one reason for relay availability issues, as most Nostr relays are operated by enthusiasts without clear financial incentives to support the operational cost.

### v2

@@ 3.2. Event Dataset
Event Collection.  We connect to each relay in our dataset to gather the posts, users and zap related events. We divide the collection process into several parallelized jobs and distributes the workload evenly across the relays, preventing any single relay from being overwhelmed.
Posts.  As described in §2, a post is a kind-1 event in Nostr, which is roughly equivalent to a tweet on Twitter/X. However, in contrast to Twitter/X, we emphasize that posts can be replicated across multiple relays (but is identified by a unique ID). Thus, we regard each unique ID as an individual post and record the relays in which it is replicated. We crawl 712 relays for the time period from 2023-07-01 to 2023-12-31, and get a total of 17,851,871 Posts.
Users.  Each unique pubkey represents an individual user. We follow this convention throughout the remainder of the paper. We identify active users who have posted at least once between 2023-07-01 and 2023-12-31 from our posts dataset. We obtain 1,558,891 unique pubkeys in total. Subsequently, we construct the user dataset by querying the relays for user profiles (kind-0 event) and the user’s following list (kind-3 event) using the obtained public keys.
Zaps.  To investigate the financial incentives of operating a relay, we collect kind-9735 events associated with the account of relay operators, which is used for the zap transactions as described in §2. We crawl the relays for the time period from 2023-07-01 to 2023-12-31, and get a total of 252,905 zap transactions, with the amount summed up to be 305,427,302 Satoshi (1 Bitcoin = 100 million Satoshi). During the measurement period, it is roughly 137,000 USD.
@@ 4. Characterizing Nostr Relays
Nostr is designed to achieve censorship resistance by distributing post replicas across multiple relays, making it difficult for attacks or failures to occur simultaneously on all relays. Thus, in this section, we focus on the properties of the Nostr relays. There are several crucial points must be considered to ensure effective censorship resistance: (i) Are the posts distributed to a sufficiently large number of relays to remain robust? (ii) How are these posts distributed across ASes and countries, and is this robust to single points of failure? (iii) How reliable are the collection of relays, hosting each individual post? (iv) What is the cost of relay operation, and is it sustainable?
@@ 4.1. Distribution of Posts & Users
Figure 1. (a) Distribution & CDF of the number of posts in a relay; (b) Distribution & CDF of the number of users in a relay.
We first evaluate the degree of decentralization within Nostr at the relay level. This gives an idea of how susceptible posts are to specific takedown attacks. To do this, we inspect the distribution of posts and users across the relays, i.e., the number of posts in the relay and the number of users using the relay. We define a user is using a relay if they have at least one post in the relay. Previous research has highlighted user-centralization and content-centralization problems in the Fediverse (Raman et al., 2019), i.e., a small number of instances host the vast majority of posts and users. While we aim to determine if these findings also apply to Nostr, we also note a vital difference: in Nostr, a user and a post can exist in multiple relays. We posit this uniqueness may result in distinctive characteristics.
Distribution of Posts & Users.  Figure 1a illustrates the distribution and cumulative distribution function (CDF) of the number of posts in a relay. We observe that the top relay hosts 73% of the posts. At first glance, this may suggest a highly centralized system. However, it is important to consider that a post can exist in multiple relays. Taking this into account, the result reveals that, although 178 relays (25%) host a significant proportion (more than 5.6%) of the posts, the sum of these proportions is far larger than 100% — on average, each post is replicated 34.6 times. This confirms that posts are extensively replicated across multiple relays. This could be viewed as a decentralized backup of posts, serving to mitigate the risk of failure or censorship in a single relay. Therefore, the result stands as a favorable signal of decentralization. We also note that this result can be biased if there is a heavy-tailed distribution, where a few posts are heavily replicated while most are not. However, we later confirm that this is not the case in §5.1, where the results are plotted in Figure 6.
Figure 1b illustrates the distribution of the number of users in a relay. We observe a similar distribution to the distribution of posts. This suggests that users are actively utilizing multiple relays for publishing (and likely retrieving) posts. Once again, this highlights Nostr’s distinctive feature of detaching users from specific servers, mitigating the risk of censorship.
Overall, the distribution patterns of posts and users demonstrates the effective decentralization at the relay level. Unlike the Fediverse, we do not find evidence for excessive user-centralization or post-centralization at the relay level in Nostr.
Figure 2. Percentage of relays, posts, and users in the top 15 regions and ASes, ranked by number of relays.
@@ 4.2. Relay Hosting
Figure 3. CDF of the percentage of relays, posts, and users in (a) the countries; (b) the ASes. The x-axis begins with a near-zero number but not 0, since relays of no posts are already excluded as described in §3.1.
We next evaluate the degree of decentralization of Nostr at the country and autonomous systems (AS) level. Spreading posts across multiple country and AS level jurisdictions is vital for delivering on Nostr’s censorship resistance goals, to prevent all administrators within a jurisdiction being forced to remove content. Like instances in the Fediverse, relays in Nostr follow a bottom-up approach, where administrators have the autonomy to choose where to host their relays. Prior work revealed that there is infrastructure-centralization in the Fediverse (Raman et al., 2019). That is, the hosting of instances and posts are centralized in a few countries and Autonomous Systems (ASes). Thus, we hypothesize that Nostr may encounter similar challenges, whereby the majority of relays fall under the remit of a small number of countries or ASes, making them vulnerable to regional or network-level failure and censorship attacks. To test this, we investigate the hosting of relays in Nostr to see if relays, posts or users are overly centralized in certain locations.
Countries.  Figure 3a displays the CDF of the percentage of relays, hosted posts, and users in different countries. Additionally, Figure 2a illustrates the percentage of relays, hosted posts, and users for the top 15 countries ranked by the number of relays. In stark contrast to the Fediverse, we observe that the distribution of relays is quite decentralized, with around 52% of countries having fewer than two relays. That said, upon closer inspection, we see that users and posts are concentrated in the top countries, such as the US, which hosts 85% of posts and 30% of users. Nonetheless, it is crucial to recall that, unlike Fediverse instances, a post and a user can belong to multiple relays, which creates backups that will persist even if some relays censor or fail. As a result, most of the top 15 countries host at least 20% of the posts, and Canada, France, Germany, and Finland host more than 40%. This distribution is actually a positive indicator of decentralization, as it suggests that a significant portion of Nostr posts and users are spread across different countries. This outcome appears to be more favorable compared to Mastodon, where a prior study revealed that 89.1% of all toots were concentrated on instances in Japan, the US, and France (Raman et al., 2019).
ASes.  Figure 3b displays the CDF of the percentage of relays, posts, and users available via different ASes. [^1] We note that some ASes listed could be Content Delivery Networks (CDNs), who do not directly host the posts themselves. We are unable to identify the specific servers behind these CDNs as currently there is no reliable approach to do so. Additionally, Figure 2b illustrates the percentage of relays, posts, and users for the top 15 ASes ranked by the number of relays. We observe a similar trend to countries, as the distribution of relays across ASes is also surprisingly decentralized: 64% of ASes host only one relay. As shown in Figure 2b, among the top 15 ASes, all of them host more than 15% of the posts. Once again, this serves as a positive indicator of decentralization and offers greater protection to AS-level censorship attacks, whereby a given AS could shutdown entire sets of relays.
Note, this distribution also contrasts significantly with the findings in the Fediverse. According to (Raman et al., 2019), toots and users in Mastodon were primarily centralized around just three major ASes, as administrators were naturally drawn to well-known and cost-effective cloud providers. One possible explanation for this disparity is that Nostr relays are typically more lightweight than Fediverse instances, requiring fewer computational resources. Consequently, administrators may find better pricing options from smaller providers rather than relying on major providers like AWS.
@@ 4.3. Relay Availability
Recall, the main design purpose behind Nostr’s greater decentralization is superior availability and resilience to censorship. Thus, we next explore the availability properties of the relays. In this context, availability refers to the ability to establish a connection with the relay. Similar to Fediverse instances, we posit that the voluntary nature of certain relay operations might lead to unique availability characteristics.
@@
Figure 4. (a) CDF of the percentage of uptime for free, paid, alive, and all relays; (b) CDF of the outage duration for free, paid, and all relays; (c) CDF of the outage duration for relays with different number of posts; (d) CDF of the count of zapped Satoshi (amount of zap), zap transactions and zappers for relays.
Relay Availability.  Based on the relay dataset described in §3, we calculate the uptime of relays from 2023-10-01 to 2023-12-31. The CDF of the uptime percentage is presented in Figure 4a. From the figure, we observe that the availability of some relays is relatively good, with 50% of the relays being available for more than 99% of the specified time period. However, there are also some relays with extremely poor availability. Approximately 20% of the relays experience downtime for more than 40% of the time. However, further investigation reveals that most of these relays are offline for extended periods and do not even reactivate by the end of our measurements. We therefore define relays that have been offline for more than a week and remain so at the end of our measurements as dead. In total, we identify 132 dead relays. When removing these dead relays from the analysis, as shown by the red line in Figure 4a, the availability of improves significantly, with 90% of the relays being operational for more than 80% of the time.
Overall, this finding confirms that relay failures are relatively common in Nostr. This broadly aligns with that observed in the Fediverse (Raman et al., 2019) (although the most significant failures can be attributed to dead relays). However, fortunately, in Nostr, users have the flexibility to switch between relays easily. As a result, the impact of relay unavailability in Nostr may not be as severe as in instance unavailability in Fediverse.
Outage Duration.  We next investigate the outage duration. For each outage, we compute its duration and plot the CDF in Figure 4b and 4c . We do not exclude dead relays in this analysis, as dead relays can be considered as relays with a long outage duration. First, we compare free and paid relays to determine if paid relays also exhibit better performance in terms of shorter outage duration. Interestingly, we find that the outage durations for paid and free relays are nearly identical, as depicted in Figure 4b. The majority of outages (56%) last for less than 15 minutes. However, there is also a long tail, with 10% of the outages lasting for more than 100 minutes.
Next, we group the relays based on the number of posts to observe if the scale of the relays correlates with the outage duration. As shown in Figure 4c, we notice that larger relays perform better, particularly in the tail end. For the group of relays with over 1 million posts, only 1% of the outages last for more than 1000 minutes, whereas the corresponding figure for the group of relays with less than 100 posts is 88%. This finding suggests that although the downtime may be similar, the outage durations for larger relays are shorter. This implies that while the availability of larger relays may not be consistently stable, it is unlikely for them to experience prolonged outages lasting several hours. On the other hand, relays that remain operational more consistently are therefore more likely to have posts successfully pushed to them. Thus, they tend to accumulate more posts.
@@ 4.4. Relay Incomes
One of the challenges within the Fediverse is the cost of running instances, coupled with a lack of incentives for doing so (Allen et al., 2023). This is also one reason why some instances have low availability, with some even being abandoned and subsequently closed. This issue also applies to Nostr, as the lack of financial income may lead to a poor availability of certain relays and the system may become centralized around well-resourced individuals (who do not require financial support). Therefore, we examine the possible incomes of the relays, and importantly, whether the income generated is sufficient to cover operational costs, and thus to keep the decentralized operation of the Nostr ecosystem. We argue that this is key to relay availability and therefore censorship resistance, as a lack of money could result in relays shutting down.
@@
Paid Relays.  One direct source of revenue is the admission fee for paid relays. This fee is a one-time cost for users to publish to the relay. We find that the admission fee is relatively low for some of the relays, with 35% charging less than 1k Satoshi (approximately \$0.45 USD). We argue that this minimal fee makes it challenging to sustain the relay’s operation. This implies that, for these relays, the primary purpose of the admission fee is to act as a barrier, allowing only genuine users while keeping spammers out, rather than generating profits. On the other hand, there are also relays that impose a higher admission fee, with 15% of relays charging more than 10k Satoshi (approximately \$45 USD). This higher fee may potentially help in covering the expenses associated with operating the relay. We explore this in the subsequent paragraphs.
@@
Zaps.  Another available revenue stream is through zaps sent by users to relay administrators. To estimate this, we analyze the zap transactions to the public keys specified in relay information. If no public key is provided in the relay information, we consider the zap income as zero. Figure 4d shows the CDF of the total zap amount (note that 52% of the transactions do not disclose the zap amount publicly), count of zap transactions, and zappers. We find that the majority (64%) of relays do not receive any zaps. Among the remaining relays, some of them do receive a relatively high number of zap transactions, with the 90th percentile at 471. However, the zap amount itself is not substantial, with the 90th percentile receiving a total of 150k Satoshi (approximately 67 USD).
Figure 5. Number of users vs. the estimated 6-month income of free and paid relays. The green line is the estimated 6-month cost of running a relay.
@@
Enough to Cover the Cost?  We finally assess whether the income generated is sufficient to cover operational costs. For this, we estimate the 6-month income (2023-07-01 to 2023-12-31) of relays. To achieve this, we first calculate the zap amount for the relays. Then, for paid relays, we add the product of the admission fee and the number of active users. For relays operational for less than 6 months, we calculate the daily average income and multiply it by 184 to estimate the income for the 6-month period. Figure 5 illustrates the estimated income vs. the number of users as a representation of the relay’s scale (See Appendix A.4 for figures with different Bitcoin prices). The green line roughly estimates the cost of operating a relay based on the cost of running a Mastodon instance of similar size, as listed in (Mastohost, 2023).
@@
It is evident that, barring a few outliers, most of the free relays with more than 10 users are unable to cover their operating costs. However, most paid relays of medium scale (with 50 – 1k users) are able to cover their operational costs. This constitutes a key threat to long-term viability. It is therefore necessary for relays to obtain additional revenue (often donated by the voluntary administrator). The trend shown in the figure is significant in terms of magnitude. Therefore, the conclusion remains valid, even though there may be some errors due to transactions not publicly disclosing the zap amount and the accuracy of estimating the operational cost based on Mastodon.
This result suggests that, akin to the Fediverse, most Nostr relays are operated by enthusiasts without clear financial incentives. The number of dead relays and the lower availability of free relays revealed in §4.3 may be a consequence of this situation. However, the observed difference in uptime between paid and free relays (§4.3) is not as stark as this financial disparity might suggest. This implies that non-monetary factors, such as hobbyist enthusiasm and the relative operational simplicity of the Nostr software, could also be drivers of reliability. Consequently, while financial unsustainability is a threat to the long-term existence of free relays, our current data does not support a strong causal link between a relay’s payment model and its day-to-day uptime.
@@
inline,color=whiteinline,color=whitetodo: inline,color=whiteTakeaways: (i) Nostr achieves greater decentralization than previous Fediverse applications, with posts replicated across multiple relays in 44 countries and 151 autonomous systems, confirming high censorship resistance. (ii) Unstable availability of relays is still a challenge for Nostr, akin to Fediverse applications. (iii) Financial sustainability might be a complicating factor in relay availability issues, as most Nostr relays are operated by enthusiasts without clear financial incentives to support the operational cost.